# MaskVCT: Masked Voice Codec Transformer for Zero-Shot Voice Conversion With Increased Controllability via Multiple Guidances

### v1

@@ 1 Introduction
Voice conversion (VC) aims to transform a source utterance to match a target’s identity while preserving the source’s linguistic content, but a key challenge is disentangling such content from pitch and timbre. Recent VC models [1, 2, 3, 4, 5, 6] used pre-trained self-supervised models’ representations as linguistic features, however, these features contains rich information to reconstruct the mel-spectrogram itself while they mentioned “disentangled” [2].
Some models are conditioned on the target’s pitch. Those pitch-conditioned models [2, 3, 5] have high controllability with frame-wise pitch, but a robust pitch predictor is needed. Meanwhile, pitch-unconditioned models [4, 6, 7] generate speech without the need for an explicit pitch input. One noticeable point is that pitch-unconditioned models can generate speech from tonal languages, such as Chinese [6], due to the entangled pitch information in pre-trained linguistic features.
@@
Many models mainly focus on converting the vocal timbre [3, 6] of the source to that of the target. Recently, zero-shot models have emerged, allowing conversion to previously unseen speaker voices without additional fine-tuning [2, 4, 5, 6]. Another trend is leveraging codec-based language models (CodecLMs). Targeting discretized tokens from neural audio codecs [8, 9, 10], CodecLMs cast generation as a Transformer-friendly classification task [11] and reconstruct via autoregressive prediction [12] or masked modeling [6], typically conditioned on a short 3-second speaker prompt.
In this work, we propose MaskVCT, a masked generative model for zero-shot voice conversion that combines temporal-quantized syllabic features with increased controllability. By using syllabic representations from SylBoost [13], MaskVCT gains flexibility in phoneme selection and accent variation within each segment. By leveraging multiple classifier-free guidances (CFGs) [14], we allow for a dynamic balance between conditions at inference time, effectively creating a highly controllable system. Unlike prior VC systems bound to a single condition, our unified model supports different conditions and inference modes in a single model, with respect to:
  • Linguistics: discrete vs. continuous representations.
  • Pitch: source-conditioned vs. target-driven contour.
  • CFG weights: balancing content, speaker, and pitch.
@@ 2.1 Linguistic Conditioning
Self-supervised speech representations provide time-aligned and fast-varying phonetic features [15] so VC models built on them tend to reproduce source-aligned speech. Table 2 shows that even pitch-unconditioned, linguistics-only-conditioned models [4, 6] retain high pitch correlation with the source, indicating pitch leakage. This motivates a coarser and pitch-stripped representation for source-pitch-agnostic conversion.
@@ 2.2 Pitch Conditioning
We map continuous pitch to a vector using sinusoidal embedding [11], making the system agnostic to pitch-extractor resolution and requiring no fine-tuning when changing extractors. Unlike the standard interleaved layout for positional embedding, we concatenate sine and cosine terms to avoid confounding multi-head attention. As the human perception of frequency is related to the log-scale, we convert pitch to the log-scale, adding an epsilon value of 1 Hz to the frequency, to prevent negative infinity with fundamental frequencies of 0 Hz. Therefore, for the frequency $f$, the $i$-th index of the pitch embedding $\mathbf{P}(f)_{i}$ is defined as follows:
$\mathbf{P}(f)_{i}=\begin{cases}\sin\left(\frac{\log(1+f)}{10000^{2\cdot i/d}}\right)&\text{if $i<d/2$}\\ \cos\left(\frac{\log(1+f)}{10000^{2\cdot(i-d/2)/d}}\right)&\text{if $i\geq d/2$}\end{cases}$ (1)
where $i$ starts from 0 to d1d-1, and $d$ is the model’s dimension. We extract all pitch contours with a rate of 50 Hz using Praat [18].
@@ 2.4 Masked Codec Language Model
The masked model is trained by reconstructing the source’s encoded acoustic tokens $\mathbf{A}_{0}\in\{1,2,\dotsc,K\}^{T\times C}$ from a masked input $\mathbf{A}_{u,c}=\mathbf{A}_{0}\odot\mathbf{M}_{u,c}$, where $K$ is the codebook vocabulary size, $T$ the number of frames, $C$ the number of the RVQ codebooks, $c\in\{0,1,\dotsc,C-1\}$ a target RVQ codebook layer, $u\in[0,1]$ a masking timestep, and $\mathbf{M}_{u,c}\in\{0,1\}^{T\times C}$ a binary mask. To be specific, for the mask $\mathbf{M}_{u,c}$, all layers below $c$ remain unmasked, all layers above $c$ are fully masked, and only layer $c$ is masked according to a masking schedule. The reconstruction process can be framed as a classification task where the model $\theta$ learns the conditional probability $p_{\theta}(\mathbf{A}_{0}|\mathbf{A}_{u,c},\mathbf{C})$, where $\mathbf{C}$ are the conditioning features. During training, we randomly sample a masking time step $u\sim\mathcal{U}[0,1]$, a target codebook layer $c\sim p(c)=1-\frac{2(c+1)}{C(C+1)}$ [6], and cosine scheduled mask $\mathbf{M}_{u,c}\sim\mathrm{Bernoulli}\left(\cos\left(\pi u/2\right)\right)$ [20]. We train our model using a simple classification loss applied exclusively to the masked tokens, defined as follows:
$\mathcal{L}_{\text{mask}}=\displaystyle\mathop{\mathbb{E}}_{t|\mathbf{M}_{u,c}[t,c]=0}\left[-\log p_{\theta}\left(\mathbf{A}_{0}[t,c]|\mathbf{A}_{u,c},\mathbf{C}\right)\right].$ (2)
During inference, we start from all-masked sequence $\mathbf{A}_{1,0}=\mathbf{A}_{N}$, and iteratively unmask to $\mathbf{A}_{0}$ over $N$ iteration, predicting logits for masked positions of codebook $c$ as $z_{\theta}(\mathbf{A}_{n}\mid\mathbf{A}_{n+1},\mathbf{C})$ and unmasking tokens at each step.
@@ 2.5 Multiple Classifier-Free Guidances
where $\omega_{\text{all}}$, $\omega_{\text{spk}}$, $\omega_{\text{ling}}$ are CFG coefficients of each conditioning, respectively. We trained a model with four different conditions, as shown in Figure 1, and all-conditioned, spk-conditioned, ling-conditioned, and null-conditioned are randomly sampled in a 6:2:2:1 ratio. For each codebook layer of audio tokens, linguistic and pitch conditions have their own mask tokens, which replace the original token to mask it.
@@ 3.1 Baselines
For our baseline, we utilized the official implementation and checkpoint of Diff-HierVC [^1] https://github.com/hayeong0/Diff-HierVC [5], FACodec [^2] https://hf.co/amphion/naturalspeech3_facodec [10, 25], MaskGCT [^3] https://hf.co/amphion/MaskGCT [6], FreeVC [^4] https://github.com/OlaWod/FreeVC [4], GenVC [^5] https://github.com/caizexin/GenVC [7]. For accent conversion, we only tested the models that achieved the best objective metrics in the main evaluation. Table 1 shows the number of parameters and training dataset hours for each models.
@@ 3.2 Training
We employ the official checkpoint of the DAC 16 kHz [8] as a codec, from which we extract 9 codebook indices to encode each audio frame. For syllabic representations, we use SylBoost 8.33 Hz [13], which has a minimum of tokens. To improve robustness to phase-related inconsistency [26], we apply PhaseAug [27] to all inputs before DAC encoder. Although SylBoost yields consistent features, self-reconstruction can leak information [2]; thus, following NANSY [2], we pitch-shift clean speech to create perturbed variants and train with a 50% property for each case.
@@ 3.3 Dataset
Our model was trained on diverse speech corpora to enhance its generalization across accents and speaking styles. Specifically, we utilized the train-clean subsets of LibriTTS‑R [29], the train subset of MLS‑en [30], VCTK [31], LibriHeavy-Large [32, 33], the clean subset of HiFi‑TTS [34], LJSpeech [35], and the speech subset of RAVDESS [36].
We utilized the test-clean subset of LibriTTS-R [29] to test the general performance of VC models. Since the models were not trained with the long length, we select the speeches that are shorter than 10 seconds and slice 3 seconds for the prompt speeches. We test the models with 511 sample pairs.

### v2

@@ 1 Introduction
Voice conversion (VC) aims to transform the identity of a source utterance to match a target speaker while preserving the original linguistic content. A fundamental challenge in this task is the effective disentanglement of linguistic content from pitch and timbre. Recent VC models [31, 11, 12, 27, 10, 37] have increasingly adopted representation from pre-trained self-supervised learning models as linguistic features. However, despite being described as “disentangled”, these features often remain “leaky”, retaining sufficient pitch and speaker information to reconstruct the original mel-spectrogram [11].
This information leakage creates a divide in model conditioning. Pitch-conditioned models [11, 12, 10] offer high controllability via frame-wise pitch guidance, but rely on accurate external pitch predictors. In contrast, pitch-unconditioned models [27, 37, 6] generate speech without explicit pitch conditioning. These models rely on the pitch information that remains entangled within linguistic features; this entanglement notably enables the generation of tonal languages like Chinese [37].
@@
Recently, zero-shot models have emerged, allowing conversion to previously unseen speaker voices without additional fine-tuning [11, 27, 10, 37]. Specifically, codec-based language models (CodecLMs) have become a dominant paradigm. By treating discretized neural audio tokens as targets [24, 20], CodecLMs formulate speech generation as a Transformer-based classification task [34], utilizing either autoregressive [36] or masked modeling [37] conditioned on short speaker prompts. Despite their popularity, however, we identify a critical limitation that zero-shot VC with prompting also struggles to discard source pitch information. This information leakage results in converted speech that retains the source’s intonation rather than adopting the target’s style.
To address these limitations of controllability and disentanglement, we propose MaskVCT, a masked generative model for zero-shot voice conversion that combines temporal-quantized syllabic features with multiple classifier-free guidances to enable enhanced controllability. By using syllabic representations from SylBoost [1], MaskVCT gains flexibility in pitch modulation, phoneme selection, and accent variation within each segment. By leveraging multiple classifier-free guidances (CFGs) [17], we allow for a dynamic balance between conditions at inference time, effectively creating a highly controllable system. Unlike prior VC models bound to a single condition, MaskVCT supports different conditions and inference modes in a single model, with respect to:
  • Linguistic representation (categorical): discrete vs. continuous.
  • Pitch (categorical): source-conditioned vs. target-driven contour.
  • CFG weights (continuous): balancing content, speaker, and pitch.
@@ 2.1 Linguistic Conditioning
Self-supervised speech representations provide time-aligned and fast-varying phonetic features [13]. Therefore, VC models built on them tend to reproduce source-aligned speech. In our preliminary study, even pitch-unconditioned and linguistics-only-conditioned models [27, 37] retain high pitch correlation with the source, indicating pitch leakage. This motivates a coarser and pitch-stripped representation for source-agnostic pitch generation.
@@ 2.2 Pitch Conditioning
We map continuous pitch to a vector using sinusoidal embedding [34], making the system agnostic to pitch-extractor resolution and requiring no fine-tuning when changing extractors. Unlike the standard interleaved layout for positional embedding, we concatenate sine and cosine terms to avoid confounding multi-head attention. As the human perception of frequency is related to the log-scale, we convert pitch to the log-scale, adding an epsilon value of 1 Hz to the frequency, to prevent negative infinity with fundamental frequencies of 0 Hz. Therefore, for the frequency $\mathbf{f}\in\mathbb{R}^{T}$, the pitch embedding $\mathbf{P}\in\mathbb{R}^{T\times d}$ is defined as follows:
$\mathbf{P}[t,i]=\begin{cases}\sin\left(\frac{\log(1+\mathbf{f}[t])}{10000^{2\cdot i/d}}\right)&\text{if $i<d/2$}\\ \cos\left(\frac{\log(1+\mathbf{f}[t])}{10000^{2\cdot(i-d/2)/d}}\right)&\text{if $i\geq d/2$}\end{cases}$ (1)
where $i$ in 0 to d1d-1, $t$ in 0 to T1T-1, and $d$ is the model’s dimension. We extract all pitch contours with a rate of 50 Hz using Praat [4].
@@ 2.4 Masked Codec Language Model
The masked model is trained by reconstructing the source’s encoded acoustic tokens $\mathbf{A}_{0}\in\{1,2,\dotsc,K\}^{T\times Q}$ from a masked input $\mathbf{A}_{u,q}=\mathbf{A}_{0}\odot\mathbf{M}_{u,q}$, where $K$ is the codebook vocabulary size, $T$ the number of frames, $Q$ the number of the RVQ codebooks, $q\in\{0,1,\dotsc,Q-1\}$ a target RVQ codebook layer, $u\in[0,1]$ a masking timestep, and $\mathbf{M}_{u,q}\in\{0,1\}^{T\times Q}$ a binary mask. To be specific, for the mask $\mathbf{M}_{u,q}$, all layers below $q$ remain unmasked, all layers above $q$ are fully masked, and only layer $q$ is masked according to a masking schedule. The reconstruction process can be framed as a classification task where the model $\theta$ learns the conditional probability $p_{\theta}(\mathbf{A}_{0}|\mathbf{A}_{u,q},\mathbf{C})$, where $\mathbf{C}$ are the conditioning features. During training, we randomly sample a masking time step $u\sim\mathcal{U}[0,1]$, a target codebook layer $q\sim p(q)=1-\frac{2(q+1)}{Q(Q+1)}$ [37], and cosine scheduled mask $\mathbf{M}_{u,q}\sim\mathrm{Bernoulli}\left(\cos\left(\pi u/2\right)\right)$ [7]. We train our model using a simple classification loss applied exclusively to the masked tokens, defined as follows:
$\mathcal{L}_{\text{mask}}=\displaystyle\mathop{\mathbb{E}}_{t|\mathbf{M}_{u,q}[t,c]=0}\left[-\log p_{\theta}\left(\mathbf{A}_{0}[t,c]|\mathbf{A}_{u,q},\mathbf{C}\right)\right].$ (2)
During inference, we start from all-masked state $\mathbf{A}_{1,0}=\mathbf{A}_{N}$, and iteratively unmask tokens over $N$ iterations, producing intermediate states $\mathbf{A}_{n}$, until reaching the fully unmasked state $\mathbf{A}_{0}$, where $n\in\{0,1,\dotsc,N\}$. Each iteration index $n$ corresponds to a codebook index $q$ according to a predefined schedule. The model estimates logits for masked positions as $\log p_{\theta}(\mathbf{A}_{n}\mid\mathbf{A}_{n+1},\mathbf{C})$.
@@ 2.5 Multiple Classifier-Free Guidances
where $\omega_{\text{all}}$, $\omega_{\text{spk}}$, $\omega_{\text{ling}}$ are CFG coefficients of each conditioning, respectively. We trained a model with four different conditions, as shown in Figure 1, and all-conditioned, spk-conditioned, ling-conditioned, and null-conditioned are randomly sampled in a 6:1:2:1 ratio. For each codebook layer of audio tokens, linguistic and pitch conditions have their own mask tokens, which replace the original token to mask it.
@@ 3.1 Baselines
For our baseline, we utilized the official implementation and checkpoint of Diff-HierVC [^1] https://github.com/hayeong0/Diff-HierVC [10], FACodec [^2] https://hf.co/amphion/naturalspeech3_facodec [20, 39], MaskGCT [^3] https://hf.co/amphion/MaskGCT [37], FreeVC [^4] https://github.com/OlaWod/FreeVC [27], GenVC [^5] https://github.com/caizexin/GenVC [6]. For accent conversion, we only tested the models that achieved the best objective metrics in the main evaluation. Table 1 shows the number of parameters and training dataset hours for each model.
@@ 3.2 Training
We employ the official checkpoint of the DAC 16 kHz [24] as a codec, from which we extract 9 codebook indices to encode each audio frame. For syllabic representations, we use SylBoost 8.33 Hz [1], which has a minimum vocabulary size. To improve robustness to phase-related inconsistency [28], we apply PhaseAug [25] to all inputs before DAC encoder. Although SylBoost yields consistent features, self-reconstruction can leak information [11]; thus, following NANSY [11], we pitch-shift clean speech to create perturbed variants and train with a 50% property for each case.
@@ 3.3 Dataset
Our model was trained on diverse corpora to enhance its generalization across accents and speaking styles. We utilized the train-clean subsets of LibriTTS‑R [23], the train subset of MLS‑en [32], VCTK [35], LibriHeavy-Large [21, 22], the clean subset of HiFi‑TTS [3], LJSpeech [18], and the speech subset of RAVDESS [29].
We utilized the test-clean subset of LibriTTS-R [23] to test the general performance of VC models. Since the models were not trained with the long length, we select the utterances that are shorter than 10 seconds and slice 3 seconds for the prompt speeches. We test the models with 511 sample pairs.
@@ 5 Conclusion
6 Acknowledgement
This work was supported by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the ARTS Program under contract D2023-2308110001. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.